OpenStudy (yisselluna):

Based on the lesson, which of the following would be the best approach for solving this system by substitution? 5x = y + 6 2x - 3y = 4 Solve the first equation for x. Solve the first equation for y. Solve the second equation for x. Solve the second equation for y.

OpenStudy (dallascowboys88):

the 2 one i got that

OpenStudy (yisselluna):

i got second one but its wrong

OpenStudy (francesca_marie):

2x -3(5x-6)=4 2x-15x + 18 = 4 -13x=-18+4 -13x = -14 x=14/13

OpenStudy (ben00):

So many people. lol

OpenStudy (francesca_marie):

then it got dead silent lol ^

OpenStudy (yisselluna):

wait i meant 1st one

OpenStudy (dallascowboys88):

yep ur right i made a typo sorry if i confused you

OpenStudy (francesca_marie):

But then again, it does say "Based on the lesson..." so i don't know how it is taught in your lesson....

OpenStudy (yisselluna):

its okay thank tho apriciate it

OpenStudy (francesca_marie):

No problem

OpenStudy (dallascowboys88):

np